OpenStudy (anonymous):

Write the expression as the sine, cosine, or tangent of an angle. cos 8x cos 2x - sin 8x sin 2x

OpenStudy (anonymous):

sum of cosines: cos(a+b)=cos(a)cos(b)-sin(a)sin(b)

OpenStudy (anonymous):

is the answer cos6x?

OpenStudy (anonymous):

I would say cos10x... the sign is opposite for the cosine

OpenStudy (anonymous):

Can you show working ?

OpenStudy (anonymous):

oh ok thanks. and yes i understand now. Can you help me with this too please? Find the exact value by using a half-angle identity. cos(pi/12)

OpenStudy (anonymous):

cos(8x)cos(2x)-sin(8x)sin(2x)=cos(8x+2x)

OpenStudy (anonymous):

the answer choices are 1/2 sqrt(1+sqrt(3) 1/2 sqrt(1-sqrt(3) 1/2 sqrt(2+sqrt(3) 1/2 sqrt(2-sqrt(3)

OpenStudy (anonymous):

\[\cos (x)=\pm \sqrt{\frac{1+\cos(2x)}{2}}\]

OpenStudy (anonymous):

is that a,b,c,or d?

OpenStudy (anonymous):

Sorry, daughter needs help getting ready for school.

OpenStudy (anonymous):

Half angle identities formula

OpenStudy (anonymous):

substitute pi/12 in for x \[\cos(\pi/12)=\pm \sqrt{\frac{1+\cos(2\frac{\pi}{12})}{2}}\]pi/12 is in the first quadrant, so it is the positive from the plus minus that we need here

OpenStudy (anonymous):

I'm sorry, I don't just give answers, but I'll work with you... what do you think the next step is?

OpenStudy (anonymous):

to simplify?

OpenStudy (anonymous):

yes... do you know the answer to the cos inside the root?

OpenStudy (anonymous):

\[\cos(2\frac{\pi}{12})=\cos(\frac{\pi}{6})=?\]you can look it up in your trig book under unit circle.

OpenStudy (anonymous):

i got 1/2 sqrt(2+sqrt(3) is that correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

You're welcome, sorry for the delay while I took my girl to school.